Please help I need to define a variable, write an inequality and solve each problem. My question is Two times the sum of a number and four is no more than three times the sum of the number and seven decreased by four.

Answer by Boreal(15235)

the number is x
2(x+4) is less than or equal to 3(x+7)-4. That says twice the sum of a number added to 4 is less than or equal to (not more than) thrice the sum of a number and seven decreased by four.
2x+8 <= 3x+21-4
2x+8 <= 3x+17
-x <= 9
x> = -9 ANSWER
Change the inequality when you multiply or divide by a negative number
try -10 should not work
(-6)2 is less than or equal to 3(-10+7)-4
-12 is less than -13. No, it does not work
----------------------------------------
-5 should work
2(-1) is less than 2(2)-4
-2 is less than 0. Yes, it works.